Entries for Bad Debt Expense Under the Direct Write-Off and Allowance Methods Seaforth International wrote off...

  1. Entries for Bad Debt Expense Under the Direct Write-Off and Allowance Methods

    Seaforth International wrote off the following accounts receivable as uncollectible for the year ending December 31:

    Customer Amount
    Kim Abel $21,550
    Lee Drake 33,925
    Jenny Green 27,565
    Mike Lamb 19,460
    Total $102,500

    The company prepared the following aging schedule for its accounts receivable on December 31:

    Aging Class (Number
    of Days Past Due)
    Receivables Balance
    on December 31
    Estimated Percent of
    Uncollectible Accounts
    0-30 days $715,000 1%
    31-60 days 310,000 2
    61-90 days 102,000 15
    91-120 days 76,000 30
    More than 120 days 97,000 60
    Total receivables $1,300,000

    a. Journalize the write-offs under the direct write-off method. If an amount box does not require an entry, leave it blank.

    b. Journalize the write-offs and the year-end adjusting entry under the allowance method, assuming that the allowance account had a beginning credit balance of $95,000 on January 1 and the company uses the analysis of receivables method. For a compound transaction, if an amount box does not require an entry, leave it blank.

    Write-off
    Adjustment

    c. How much higher (lower) would Seaforth International's net income have been under the allowance method than under the direct write-off method?
      $

Check My Work2 more Check My Work uses remaining.
Answer #1

Estimated uncollectible amount = (715000*1%+310000*2%+102000*15%+76000*30%+97000*60%) = 109650

a) Journal entry

Date account and explanation debit credit
bad debt expense 102500
Allowance for doubtful accounts 102500
(To record write off)

b) Journal entry

Date account and explanation debit credit
Allowance for doubtful accounts 102500
Account receivable 102500
(To record write off)
Bad debt expense (109650+7500) 117150
Allowance for doubtful accounts 117150
(To record bad debt expense)

c) Net income will be lower under allowance method by (117150-102500) = $14650